Assume you purchase a $225,000 home. What would your monthly payment be with a 4% annual rate on a 30 year mortgage with a 15% down payment? Use excel.
Answer #1

We use the excel function to find the monthly payment.

Monthly payment = $913.0568

B D A 1 Home price 2 Down payment 3 Loan amount C $225,000 $33,750 $191,250 15% 5 Interest rate 6 Monthly interest rate 7 Mor

Screenshot with formulas

A в 1 Home price 2 Down payment 3 Loan amount 0.15 225000 =C1*B2 =C1-C2 5 Interest rate 6 Monthly interest rate 7 Mortgage te

Can you please upvote? Thank You :-)